# Why do Learning Rates Transfer? Reconciling Optimization and Scaling Limits for Deep Learning

### v1

@@ 2.2 Conditions for hyperparameter transfer
The empirical success of hyperparameter transfer crucially relies on the following two observations/desiderata.
@@
    The optimal learning rate is consistent across widths/depths (hyperparameter transfer).
@@
From these observations, we conclude that transfer does not necessarily imply matching loss dynamics [^2] In our toy setting of Section 4, loss dynamics at different widths are very similar since the hypothesis class is fixed across widths. This is not the case for our deep learning experiments where the hypothesis class is larger (and thus lower loss can be achieved) as the model size increases.. In this paper, we argue that there are instead properties of the loss landscape along the SGD trajectory that effectively transfer in a largely width/depth independent way under the right scaling of the architecture. Thus, the optimal learning rate is preserved across width/depth throughout a sizeable fraction of the training trajectory.
2.3 Optimization: Optimal Learning Rate and Edge of Stability
We analyze the sharpness $\lambda_{t}$, defined as the largest eigenvalue of the Hessian: $\nabla^{2}_{\theta}\mathcal{L}(\theta_{t}):=\sum_{\mu}\nabla^{2}_{\theta}\mathcal{L}(f_{t}(x_{\mu}))\in\mathbb{R}^{P\times P}$, as $\theta_{t}$ evolves with gradient descent. There are an abundance of works in the literature that have studied the structure of the Hessian in deep models (Singh et al., 2021, Martens, 2020, Orvieto et al., 2021). The sharpness has been linked to various aspects of training speed and stability, and often provides a theoretically principled upper bound on the maximum step size allowed. For instance, for a quadratic objective, $\lambda_{t}=\lambda$ is constant and gradient descent would diverge if the learning rate $\eta>2/\lambda$, and training speed is maximized for $\eta=1/\lambda$ (LeCun et al. (2002), page 28). Beyond this classical example, the descent lemma (Nesterov, 2013) states [^3] The result is generally stated using the gradient Lipschitz constant. This corresponds to the top Hessian eigenvalue on twice continuously differentiable losses. that for a general non-convex loss $\mathcal{L}$,
$\mathcal{L}(\theta_{t+1})-\mathcal{L}(\theta_{t})\leq-\left(\eta-\frac{\eta^{2}\lambda_{t}}{2}\right)\|\nabla\mathcal{L}(\theta_{t})\|^{2}.$
The lemma predicts that $\mathcal{L}(\theta_{t+1})-\mathcal{L}(\theta_{t})$ if $\eta\leq\frac{2}{\beta}$ where $\beta:=\sup_{\theta}\|\nabla^{2}L(\theta)\|_{2}$, where $\|\nabla^{2}L(\theta)\|_{2}$ is the sharpness at $\theta$. When it comes to deep neural networks, $\lambda_{t}$ is generally observed to increase during training (progressive sharpening): in the early phase of training it increases Jastrzkebski et al. (2018), Jastrzebski et al. (2020) (depending on the learning rate and the batch size) and then it decreases close to convergence Sagun et al. (2016). Under full batch gradient descent training, the sharpness consistently rises above the Edge of Stability (EoS) threshold of $2/\eta$.
Figure 3: Depth-extension on ResNet (Eq. 1) with convolutional layers and ReLU activation trained on CIFAR-10 for 80 epochs (with data augmentation). The bottom plot shows the sharpness for the first 40 epochs of training. The width is fixed at 128.
() Random Features
() $\mu$P
() NTP
3 Why Do Learning Rates Transfer?
Motivated by the intrinsic dependencies between the sharpness $\lambda$ and optimal learning rate, we study the time evolution of $\lambda$ as a function of width and observe the following:
Observation 1: in $\mu$P (and $\mu$P-$\sqrt{L}$), optimization happens along trajectories of $N,L-$independent sharpness $\lambda$, while for NTP the sharpness decreases in width.
We remark that under the observation that the sharpness $\lambda$ controls the maximum step size allowed, having $N,L$-independent trajectories provides supporting evidence for learning rate transfer, as the optimal learning rate will also be $N,L$-independent.
@@ 3 Why Do Learning Rates Transfer?
In Fig. 1 we train a two-layer convolutional network under the $\mu$P and NTP scalings with cross entropy loss, while keeping track of the sharpness at fixed gradient step intervals. The top row shows the dynamics of $\lambda$. Notice how the sharpness’ behaviour is qualitatively different in the two parameterizations: in $\mu$P it reaches a width-independent value which is close to the stability threshold of $2/\eta_{0}$. On the other hand, in NTP we observe a progressive diminishing of the sharpness with width, as previously observed for Mean-Square-Error loss by Cohen et al. (2021).
@@
We then study the effect of depth under the $\mu$P-$\sqrt{L}$ model of Eq. 1. In Fig. 3, we show that the sharpness’ dynamics are also extremely consistent across depth, although progressively diminishing from the EoS threshold. This suggests that EoS is not necessary for the learning rate to transfer, but the consistency of the sharpness’ dynamics is.
$\boldsymbol{N,L}$-independence of the sharpness and loss curves.
In order to have meaningful learning rate transfer (as described in the desiderata in Sec. 2.2), it is necessary that the dynamics of the sharpness stay $N,L$-independent along large parts of the training trajectory, while the higher capacity models (larger width) are able to fit data faster. In other words, $\lambda$ must converge faster than the predictor $f(x)$ with respect to the scaling quantity (width or depth) at any time step. Indeed, this is what we observe:
Observation 2: in $\mu$P (and $\mu$P-$\sqrt{L}$), the $N,L$-independent sharpness is sustained for a long period of training time, while the loss curves progressively depart from each other. In NTP, the sharpness shows inconsistent dynamics from early stages of training.
While the observation of a consistent sharpness at a large time scale can already be observed in Fig. 4 (bottom row), we now directly compare loss curves with sharpness dynamics under $\mu$P and NTP. In Fig. 2, we study the early training dynamics (first 200 gradient steps) of a 3-layer convolutional network under $\mu$P and NTP. Notice how for $\mu$P: (1) the training curves progressively depart from each other — with wider being better — while (2) the sharpness stays at its EoS value of $2/\eta_{0}$ for a sustained period of the training trajectory. Also, notice $\lambda$ reaches the EoS value very consistently in width (only the width 8 model shows slower convergence). On the other hand, under the NTP parametrization, the sharpness decreases in width, and its progressive sharpening phase converges to EoS at different speeds, with shallower models converging faster.
The observation that sharpness dynamics exhibit greater consistency compared to loss dynamics suggests that under $\mu$P scaling, although models with larger capacity fit the data faster, the paths taken by various models through the optimization landscape show a surprisingly uniform curvature.
@@ 3.1 Feature Learning and Progressive Sharpening
Observations 1 and 2 provide empirical evidence as to why — from an optimization perspective — $\mu$P exhibits learning rate transfer. Along the lines of Yang et al. (2023) and Bordelon et al. (2023), but with the additional insights from our experiment, we conjecture that under NTP the progressive lack of feature learning and the convergence of the model to its linearization prevents the progressive sharpening needed to reach a $N,L$-independent $\lambda$. By increasing the learning rate $\eta_{0}$, however, the convergence speed to the infinite width limit slows down Lee et al. (2019), and thus larger width networks effectively learn features. This causes the incremental shift of optimal learning rate, as we observe in Figures 1 and 4. This brings us to the conjecture that a feature learning limit is necessary for hyperparameter transfer in deep neural networks.
@@
Indeed, the connection between feature learning and progressive sharpening becomes evident when studying the relationship between the sharpness and the largest eigenvalue of the NTK matrix. Following the Gauss-Newton decomposition Martens (2016; 2020), the Hessian can be decomposed as a sum of two matrices $H=\mathcal{G}+\mathcal{R}$, where $\mathcal{G}$ is the Gauss-Newton (GN) matrix and $\mathcal{R}$ depends on the Hessian of the model. For MSE loss,
$\mathcal{G}=\sum_{i=1}^{|\mathcal{D}|}\nabla_{\theta}f(x_{i})\nabla_{\theta}f(x_{i})^{\top}=K^{\top}K.$
where $K\in\mathbb{R}^{|\mathcal{D}|\times P}$ is a matrix where each row is $\nabla_{\theta}f(x_{i})$ (i.e. the Jacobian of $f(x_{i})$), and:
$\mathcal{R}=\sum_{i=1}^{|\mathcal{D}|}\nabla_{\theta}^{2}f(x_{i})(y_{i}-f(x_{i})),$
@@
where $y_{i}\in\mathbb{R}$ is the label. One can readily see that the NTK matrix in Eq. 4 can be written as $\Theta(f_{\theta})=KK^{\top}$. Thus, we can immediately see [^4] Using a simple SVD. that the NTK and $\mathcal{G}$ share the same nonzero eigenvalues, and in particular the same sharpness ($\lambda(\Theta)=\lambda(\mathcal{G})$). In Figure 5, we show that under $\mu$P the sharpness evolution is dominated by the $\mathcal{G}$ consistently across different widths, while for NTP the sharpness evolution slows down when increasing the width. Since in the limit the NTK matrix is fixed for NTP, while it evolves with time for $\mu$P, these results provide further supporting evidence for the role of feature learning in the evolution of the hessian, and thus in hyperparameter transfer. While this argument strictly holds for MSE loss, it can be generalized to any twice differentiable loss function, albeit with some caveats. In Appendix D, we generalize the setting, analyze the cross-entropy loss and perform validating experiments, confirming the conclusions drawn here.
To test this conjecture further, we compare the NTP and $\mu$P parametrizations to a random feature model, i.e. a model where we all the weights of the intermediate layers are frozen to their value at initialization, and only the final readout layer is trained. Crucially, this model does not learn features by construction for any learning rate and any width. The results are shown in Fig. 4. Notice how the transfer in the random feature model is achieved only at very large widths compared to $\mu$P. However, the transfer is better than in NTP. This is in line with our claim, as under a random feature model increasing the learning rate does not induce more feature learning, as is the case in NTP.
In Section 4 we revisit the above claims more precisely in a simplified setting, providing further intuition on the sharpness dynamics and learning rate transfer.
@@
Figure 5: Evolution of the top eigenvalues of the Hessian components $\mathcal{G}$ and $\mathcal{R}$ for a two-layer linear network trained on random data under MSE loss. The vector field characterizes the evolution during training for a fixed learning rate. Left: $\mu$P. Note that all the curves follow the same trajectory and begin with a sharpening phase, after which the residual continues to decrease. Right: NTP. For wider networks the sharpening phase reduces, since the network is approaching the limit where the NTK is fixed to its value at initialization.
3.2 Other experiments
Large scale experiments.
@@ 3.2 Other experiments
In App. E.3, (Figures 11, 12, 13, 14, 15) we perform more experiments to validate the connection between the consistency of the sharpness’ dynamics and learning rate transfers across datasets (Tiny-ImageNet, WikiText), architectures (ViT, GPT-2 Radford et al. (2019)), and optimizers (Adam Kingma and Ba (2014)). We find these results to be consistent with those in the main text.
End of training dynamics.
@@
In App. E.1 (Fig. 8), we study the width dependence of the sharpness at the late phase of training. It is well-known that for cross-entropy loss, a phase transition happens where the sharpness starts to decrease Cohen et al. (2021). We found that even for $\mu$P this transition point is width-dependent, with a consequent slight shift in optimal learning rates during this late phase. Again, these results are in line with our results that width-independent sharpness facilitates transfer.
Batch size ablation.
@@
We repeat the experiment in Fig. 1 with increasing batch size, observing that the width-independent threshold is reached across all the tested batch sizes, thus not affecting learning rate transfers. For larger batches, a close-to-EoS threshold is reached across more learning rates. Results are summarized in Fig. 9 and 10 in App. E.2.
@@ 4 Case study: Two-Layer Linear Network
In this section, we revisit and validate our intuition and empirical findings in Sec. 3 through the lens of a two-layer neural network with linear activations and $L2$ loss. Our purpose is to characterize the dynamics of $\mu P$ and NTP at the edge of stability through the lens of a simple example that shares a similar phenomenology with the more complex scenarios observed in the last section (see Fig. 6, App. C). Our setting is similar to the one leading to the insightful analysis of Edge of Stability in Kalra et al. (2023), Song and Yun (2023). Compared to these works, we do not limit the analysis to a single datapoint or to vanishing targets  [^5] If our dataset has cardinality $1$, then the NTK is a scalar. If targets vanish, for 2-layer linear networks with $L2$ loss, NTP and $\mu P$ induce the same loss on parameters ($\gamma$ cancels)..
@@
Consider a dataset of $|\mathcal{D}|$ datapoints in $D$ dimensions $X\in\mathbb{R}^{|\mathcal{D}|\times D}$ ($|\mathcal{D}|\geq D$), and labels generated through a latent ground-truth vector $w_{*}\in\mathbb{R}^{D}$, that is $Y=Xw_{*}$. The neural network we use here is parametrized by weights $W^{0}\in\mathbb{R}^{D\times N}$, $W^{1}\in\mathbb{R}^{N\times 1}$, where $N$ is the width. To simplify the notation in our setting, we name $E:=W^{0}$ and $V:=W^{1}$: $f(X)=\frac{1}{\gamma\sqrt{ND}}XEV$, $\mathcal{L}(E,V)=\frac{1}{2}\|f(X)-Y\|^{2}$. We initialize each entry of $E,V$ i.i.d. Gaussian with mean zero and variance $1$. Recall that $\gamma_{\text{\tiny NTP}}=1$, $\gamma_{\text{\tiny$\mu P$ }}=\sqrt{N}$. We train with gradient descent (GD) with a learning rate $\eta=\eta_{0}\gamma^{2}$. Empirically, we observed (Fig. 6, App. C) that picking $|\mathcal{D}|=D$ and data $X=I_{D}$ ($I_{D}$ is the $D\times D$ identity matrix) is sufficient to track most of the crucial features of $\mu P$/ NTP explored in this paper, except the “wider is better” effect which here is less apparent due to the simple hypothesis class. Note that in this setting $|\mathcal{D}|$ cancels and $XE=E$. Thus, $E$ can be interpreted as the matrix of features after the first layer. The loss function reduces to
@@
Hessian top eigenvalue, GN and NTK.
Note that using a learning rate $\eta_{0}\gamma^{2}$ when optimizing $\mathcal{L}$ is equivalent to using a learning rate $\eta_{0}$ when optimizing $\gamma^{2}\mathcal{L}$. We now derive closed-form expressions for metrics on $\nabla^{2}\mathcal{L}$. Define
@@
The following bound leverages a Gauss-Newton decomposition and leads analytical insights supporting Sec. 3.1.

### v2

@@ 4 Super Consistency and Learning Rate Transfer
Conditions for hyperparameter transfer. The empirical success of hyperparameter transfer crucially relies on the following two observations, constituing a “theoretical puzzle” Yang et al. (2023).
@@
    The optimal learning rate is preserved across widths/depths, indicating very fast convergence with respect to the scaling quantity.
@@
In this section, we study the role of Super Consistency in learning rate transfer. We focus on the dynamics of sharpness $\lambda_{\max}$ across training, due to its well-established connection to optimization theory and step size selection, as well as better computational tractability than the full Hessian spectrum. We provide extensive studies of other relevant spectral quantities (i.e. Hessian and NTK eigenvalues) in Appendix G.
Observation: in $\mu$P (and Depth-$\mu$P), the sharpness $\lambda$ is super consistent along the training trajectory, while for NTP the sharpness decreases in width. This correlates with presence/absence of hyperparameter transfer.
@@
In Fig. 1 we train a two-layer convolutional network under the $\mu$P and NTP scalings with cross entropy loss, while keeping track of the sharpness at fixed gradient step intervals. The top row shows the dynamics of $\lambda$. Notice how the sharpness’ behaviour is qualitatively different in the two parameterizations: in $\mu$P it reaches a width-independent value which is close to the EoS threshold of $2/\eta_{0}$. On the other hand, in NTP we observe a progressive diminishing of the sharpness with width, as previously observed for Mean-Square-Error loss by Cohen et al. (2021).
@@
We then study the effect of depth under the Depth-$\mu$P model of Eq. 1. In Fig. 4 (left), we show that the sharpness’ dynamics are also super consistent across depth, although progressively diminishing from the EoS threshold. This suggests that EoS is not necessary for the learning rate to transfer, but the consistency of the sharpness dynamics is.
Other feature learning parameterizations. Finally, we study the effect of other feature parameterizations that do not exhibit learning rate transfer. In particular, we study the Depth-$\mu$P scaling of the residual branches in residual networks with multiple layers per residual branch - denoted by $k$ (i.e. each branch has multiple weight matrices and non linearities). A typical example is the Transformer architecture, which has multiple layers per block in both the attention and fully connected blocks. This parameterization, although it learns features in the infinite depth limit, it is lazy within each residual branch Yang et al. (2023), Bordelon et al. (2023). The results are in Fig. 4. Notice how the sharpness dynamics are not super consistent, in that they accumulate finite-size effects over time. We study other parametrizations, including those without a stable limit in Appendix B, showing compatible results with those presented here. The observation that sharpness dynamics exhibit greater consistency compared to loss dynamics suggests that under $\mu$P scaling, although models with larger capacity fit the data faster, the paths taken by various models through the optimization landscape show a surprisingly uniform curvature.
() Depth-$\mu$P ConvNet $k=1$
() Depth-$\mu$P ConvNet $k=2$
() Depth-$\mu$P ViT
Figure 4: Depth-$\mu$P extensions with top row showing transfer plots and bottom row the sharpness evolution. (a) ConvNets with $1$ layer per block exhibit both hyperparameter transfer and sharpness Super Consistency. (b) ConvNets with $2$ layers per block. The model has a lazy behavior within each block, and no transfer. The sharpness starts accumulating finite-size effects during training, violating Super Consistency. (c) ViTs also have $k>2$ blocks per layer by design, and thus have a similar behaviour. Details: (a), (b) are trained with SGD, with widths $128$ and $32$ respectively; (c) is trained with Adam, with the learning rate scaled by $1/\sqrt{L}$ Yang et al. (2023). See Fig. 22 for convergence rates.
@@ 4.1 Feature Learning and Progressive Sharpening
We now study the effect of feature learning in the sharpness dynamics. Following the Gauss-Newton decomposition Martens (2016; 2020), the Hessian can be decomposed as a sum of two matrices $H=\mathcal{G}+\mathcal{R}$, where $\mathcal{G}$ is the Gauss-Newton (GN) matrix and $\mathcal{R}$ depends on the Hessian of the model. For MSE loss,
$\mathcal{G}=\sum_{i=1}^{|\mathcal{D}|}\nabla_{\theta}f(x_{i})\nabla_{\theta}f(x_{i})^{\top}=K^{\top}K\quad\text{ and }\quad\mathcal{R}=\sum_{i=1}^{|\mathcal{D}|}\nabla_{\theta}^{2}f(x_{i})(y_{i}-f(x_{i})),$
@@
where $K\in\mathbb{R}^{|\mathcal{D}|\times P}$ is a matrix where each row is $\nabla_{\theta}f(x_{i})$ (i.e. the Jacobian of $f(x_{i})$), and $y_{i}\in\mathbb{R}$ is the label. One can readily see that the NTK matrix can be written as $\Theta(f_{\theta})=KK^{\top}$, thus the NTK and $\mathcal{G}$ share the same nonzero eigenvalues. In Figure 5, we show that under $\mu$P the sharpness evolution is dominated by the $\mathcal{G}$ matrix consistently across different widths, while for NTP the sharpness evolution slows down when increasing the width. Since in the limit the NTK matrix is fixed for NTP, while it evolves with time for $\mu$P, these results provide further supporting evidence for the role of feature learning in the evolution of the hessian. While this argument strictly holds for MSE loss, it can be generalized to any twice differentiable loss function, albeit with some caveats. In Appendix D, we generalize the setting, analyze the cross-entropy loss and perform validating experiments, confirming the conclusions drawn here. Finally, in Appendix H, we show that our results remains valid in a random feature model, where the NTK matrix is fixed at initialization at any finite width. In Section 5 we revisit the above claims more precisely in a simplified setting, providing further intuition on the sharpness dynamics and learning rate transfer.
@@
Figure 5: Evolution of the top eigenvalues of the Hessian components $\mathcal{G}$ and $\mathcal{R}$ for a two-layer linear network trained on random data under MSE loss. The vector field characterizes the evolution during training for a fixed learning rate. Top: $\mu$P. Note how $\mathcal{G}$ drives the initial change super consistently. Bottom: NTP. For wider networks the sharpening phase reduces, since the network is approaching the limit where the NTK is fixed to its value at initialization.
@@
Large scale experiments. In App. F, we perform more experiments to validate the connection between the consistency of the sharpness’ dynamics and learning rate transfers across datasets (Tiny-ImageNet, Wikitext), architectures (ViT, GPT-2 Radford et al. (2019)), and optimizers (Adam Kingma and Ba (2014) and AdamW Loshchilov (2017)). We find these results to be consistent with those in the main text.
@@
End of training dynamics. In App. E.1 (Fig. 12), we study the width dependence of the sharpness at the late phase of training. It is well-known that for cross-entropy loss, a phase transition happens where the sharpness starts to decrease Cohen et al. (2021). We found that even for $\mu$P this transition point is width-dependent, with a consequent slight shift in optimal learning rates during this late phase. Again, these results are in line with our results that super consistent sharpness facilitates transfer.
@@
Batch size ablation. We repeat the experiment in Fig. 1 with increasing batch size, observing that the threshold is reached across all the tested batch sizes, thus not affecting learning rate transfers. For larger batches, a close-to-EoS threshold is reached across more learning rates. Results are summarized in Fig. 13 and 14 in App. E.
@@ 5 Case study: Two-Layer Linear Network
We now revisit and validate our intuition and empirical findings in Sec. 4 through the lens of a two-layer neural network with linear activations and $L2$ loss. Our purpose is to characterize the dynamics of $\mu P$ and NTP at the edge of stability through the lens of a simple example that shares a similar phenomenology with the more complex scenarios observed in the last section (see Fig. 10, App. C). In particular, the theory justifies the preconditioned Hessian $\gamma^{2}H\propto NH$ as the right object of study when it comes to the sharpness computations (see Prop. 5.3). Also, it provides an intuition to the width-independent evolution of the sharpness. Our setting is similar to the one leading to the insightful analysis of EoS in (Kalra et al., 2023, Song and Yun, 2023). Compared to these works, we do not limit the analysis to a single datapoint or to vanishing targets  [^1] If our dataset has cardinality $1$, then the NTK is a scalar. If targets vanish, for 2-layer linear networks with $L2$ loss, NTP and $\mu P$ induce the same loss on the parameters ($\gamma$ cancels)..
@@
Consider a dataset of $|\mathcal{D}|$ datapoints in $D$ dimensions $X\in\mathbb{R}^{|\mathcal{D}|\times D}$ ($|\mathcal{D}|\geq D$), and labels generated through a latent ground-truth vector $w_{*}\in\mathbb{R}^{D}$, that is $Y=Xw_{*}$. The neural network we use here is parametrized by weights $W^{0}\in\mathbb{R}^{D\times N}$, $W^{1}\in\mathbb{R}^{N\times 1}$, where $N$ is the width. To simplify the notation in our setting, we name $E:=W^{0}$ and $V:=W^{1}$: $f(X)=\frac{1}{\gamma\sqrt{ND}}XEV$, $\mathcal{L}(E,V)=\frac{1}{2}\|f(X)-Y\|^{2}$. We initialize each entry of $E,V$ i.i.d. Gaussian with mean zero and variance $1$. Recall that $\gamma_{\text{\tiny NTP}}=1$, $\gamma_{\text{\tiny$\mu P$ }}=\sqrt{N}$. We train with gradient descent (GD) with a learning rate $\eta=\eta_{0}\gamma^{2}$. Empirically, we observed (Fig. 10, App. C) that picking $|\mathcal{D}|=D$ and data $X=I_{D}$ ($I_{D}$ is the $D\times D$ identity matrix) is sufficient to track most of the crucial features of $\mu P$/ NTP explored in this paper, except the “wider is better” effect which here is less apparent due to the simple hypothesis class. The loss function reduces to:
@@
Finally, we reparametrize the model by defining:
@@
Note that using a learning rate $\eta_{0}\gamma^{2}$ when optimizing $\mathcal{L}$ is equivalent to using a learning rate $\eta_{0}$ when optimizing $\gamma^{2}\mathcal{L}$. Next, we characterize how $e,v$ evolve through time, and give conclusions for $\mu$P.
5.1 Dynamics and Edge of Stability in Latent Space
We now show that at any value of the width $N$, under GD on the original network parameters $(E,V)$, the dynamics of $w,e,v$, can be described completely through a self-contained dynamical system in $(1+D+D^{2})$ dimensions. This property is surprising because the original dynamical system described by GD on the variables $E,V$ lives in $N(D+1)$ dimensions. Concretely, this means we can study the Hessian dynamics at different network widths in the same space.
Theorem 5.1 (Evolution Laws).
Let $(E,V)$ evolve with GD at stepsize $\eta=\eta_{0}\gamma^{2}$ on the loss of Eq. 4. The evolution of $(w,e,v)$ is completely described by the following self-contained equation: let the + denote updated quantities,
While the system above describes the evolution laws $(w_{k},e_{k},v_{k})\to(w_{k+1},e_{k+1},v_{k+1})$, the dynamics are influenced also by initialization. In Prop. C.1 in the Appendix, we show that the only dependence in width in the evolutions laws are in the initial conditions.
Last, by analyzing the stability of the dynamical system in Theorem 5.1, we can characterize the edge of stability using tools from dynamical systems (Ott, 2002). First of all, we need the following Lemma, which implies that at the minimizer ($w=w_{*}$), the Hessian has the same non-zero eigenvalues as the NTK $\Theta$, which only depends on $e$ and $v$.